OpenStudy (anonymous):

How many cones of radius 1 m and height 1 m have total volume equal to the volume of a sphere with radius 1 m? a. 3 c. 5 b. 4 d. 6

OpenStudy (anonymous):

Solve for x x * 1/3 pi r^2 h = 4/3 pi r^3 With r=1 and h=1

OpenStudy (anonymous):

@Nathan28 What do you get?